OpenStudy (anonymous):

What are the zero(s) of the function f(x) = the quantity of 5 x squared minus 25 x, all over x?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (campbell_st):

\[\frac{5x^2 - 25x}{x}\] what common factor do you see..?

OpenStudy (anonymous):

5x

OpenStudy (campbell_st):

ok... so I'll do that \[\frac{5x(x - 5)}{x}\] what common factor do you see in the numerator and denominator..?

OpenStudy (anonymous):

x

OpenStudy (campbell_st):

ok so that leaves 5(x -5) = 0 so what value of x makes the equation true..?

OpenStudy (anonymous):

5, 0

OpenStudy (campbell_st):

only 5 and when x = 0 in the original equation you have a zero denominator, which is undefined so at x = 0 you have a point of discontinuity... so the solution is x = 5 which comes from the simplified version of the original equation. hope it makes sense.

OpenStudy (anonymous):

It does! Thank-you:)